# Vinicius G. Goecks DEVCOM Army Research Laboratory Aberdeen Proving Ground, Maryland, USA

### 01 Feb 2024 v1

@@ Abstract
The development of Courses of Action (COAs) in military operations is traditionally a time-consuming and intricate process. Addressing this challenge, this study introduces COA-GPT, a novel algorithm employing Large Language Models (LLMs) for rapid and efficient generation of valid COAs. COA-GPT incorporates military doctrine and domain expertise to LLMs through in-context learning, allowing commanders to input mission information – in both text and image formats – and receive strategically aligned COAs for review and approval. Uniquely, COA-GPT not only accelerates COA development, producing initial COAs within seconds, but also facilitates real-time refinement based on commander feedback. This work evaluates COA-GPT in a military-relevant scenario within a militarized version of the StarCraft II game, comparing its performance against state-of-the-art reinforcement learning algorithms. Our results demonstrate COA-GPT’s superiority in generating strategically sound COAs more swiftly, with added benefits of enhanced adaptability and alignment with commander intentions. COA-GPT’s capability to rapidly adapt and update COAs during missions presents a transformative potential for military planning, particularly in addressing planning discrepancies and capitalizing on emergent windows of opportunities.
@@ I Introduction
The development of Courses of Action (COAs), fundamental to the decision-making process, exemplifies these challenges. COA development has traditionally been a meticulous and time-consuming process, heavily dependent on the experience and expertise of military personnel. However, the rapid evolution of modern warfare demands more efficient methods for developing and analyzing COAs.
@@
COA-GPT leverages LLMs to swiftly develop valid COAs, integrating military doctrine and domain expertise directly into the system’s initial prompts. Commanders can input mission specifics and force descriptions – in both text and image formats, receiving multiple, strategically aligned COAs in a matter of seconds. This breakthrough significantly reduces COA development time, fostering a more agile decision-making process. It enables commanders to rapidly iterate and refine COAs, keeping pace with the ever-changing conditions of the battlefield and maintaining a strategic edge over adversaries.
@@
  • COA-GPT, a novel framework leveraging LLMs for accelerated development and analysis of COAs, integrating military doctrine and domain expertise effectively.
@@ II Military Relevance
The current C2 processes in military operations are predominantly linear and involve numerous sequential steps, which can be cumbersome and slow in the fast-paced environment of modern warfare [3, 2]. The advent of AI in military planning, particularly the COA-GPT system, presents an opportunity to radically streamline these processes. COA-GPT’s ability to quickly generate, analyze, and refine multiple COAs can transform traditional C2 operations, combining and executing multiple planning steps concurrently to optimize efficiency.
With COA-GPT, the MDMP can be significantly enhanced. By enabling simultaneous development and analysis of COAs, COA-GPT facilitates quicker decision-making and more effective control of operations. This leads to faster generation of actionable intelligence, thereby enabling commanders to make more informed decisions swiftly. The integration of AI-based tools like COA-GPT in C2 processes can reduce the physical footprint of military operations, diminishing the need for extensive personnel and logistics support, supporting the vision of distributed commander posts.
The proposed COA-GPT system has the potential to revolutionize C2 operations by enabling rapid generation, analysis, and comparison of multiple COA alternatives. This system can integrate with war gaming simulators and process data from battlefield sensors in real-time, offering commanders the ability to quickly adapt to the dynamic battlefield environment. This vision aligns with the ideas presented in [5], which emphasize the integration of COA analysis during the development phase for rapid optimization and comparison of multiple COAs. However, COA-GPT extends this concept by leveraging the combined strengths of human expertise and AI, fostering creativity, exploration, and informed decision-making across the C2 process.
COA-GPT is designed to present COAs to personnel in an intuitive manner, using language and annotated maps, facilitating the analysis of COA alternatives and allowing C2 personnel to adapt AI-generated COAs using their domain expertise and situational awareness, and ultimately making informed decisions on final COA selections.
In practical terms, C2 personnel will dynamically interact with COA-GPT to modify proposed COAs. They can specify objectives, input data in both text and image formats, issue corrections to fine-tune plans, and set constraints to avoid undesirable actions. In response to unforeseen changes, feedback can be relayed to COA-GPT to adapt the plan to new risks and opportunities. Personnel can also adjust COA selection criteria to balance multiple objectives, such as prioritizing control over an enemy asset or minimizing danger to forces. This interactive process ensures that the final COA selected is aligned with the commander’s strategic intent and situational requirements.
@@ III-A Planning with Large Language Models
The integration of Large Language Models (LLMs) in plan of action development is revolutionizing various sectors, including disaster management and military operations. In the realm of disaster response, the DisasterResponseGPT [15] algorithm stands out. This algorithm leverages the capabilities of LLMs to quickly generate viable action plans, integrating essential disaster response guidelines within the initial prompt. Similar to the proposed COA-GPT, DisasterResponseGPT facilitates the rapid generation of multiple plans from user-inputted scenarios, significantly reducing response times and enabling real-time plan adjustments based on ongoing feedback. This methodology parallels the objectives of COA-GPT, however it hasn’t been shown to incorporate COA analysis and spatial information during COA generation.
Similar to military operations, disaster response demands rapid, informed decision-making under severe time constraints and high-pressure conditions [16, 17, 18]. Traditionally, action plan development in such contexts has been a laborious process, heavily reliant on the experience and expertise of the personnel involved. Given the stakes involved, where delays can result in loss of life, there is a critical need for more efficient and reliable plan development methodologies [19, 20, 16, 21]. COA-GPT addresses similar challenges in military operations, offering a swift and adaptive approach to COA generation and analysis.
The recent studies by Ahn et al. (2022) [12] and Mees et al. (2023) [14] expand the scope of LLM applications to include robotic and visual affordances, illustrating how LLMs can be grounded in real-world contexts. These developments are crucial in bridging the gap between high-level semantic knowledge and practical, executable actions in dynamic environments. Likewise, COA-GPT seeks to apply LLMs in a military context, translating abstract strategic concepts and battlefield information into concrete, actionable military plans.
Continuing to show the potential of LLMs for planning, the introduction of Voyager by Wang et al. (2023) [22], an LLM-powered embodied lifelong learning agent in Minecraft, represents a groundbreaking step in autonomous, continual learning and skill acquisition in complex, ever-changing environments. Similarly, the STEVE-1 [23] model showcases the potential of LLMs in guiding agents to follow complex instructions in virtual settings, leveraging advancements in text-conditioned image generation and decision-making algorithms. The adaptability and learning capabilities demonstrated in these models are qualities that COA-GPT harness for generating and refining in real-time COAs for military scenarios.
@@ III-B AI for Military Planning and Operations
The application of AI and LLMs in military planning and operations is a field of growing interest and significant potential. The First-Year Report of ARL Director’s Strategic Initiative, focusing on Artificial Intelligence for Command and Control (C2) of Multi-Domain Operations [24], exemplifies this trend. The report discusses ongoing research into whether deep reinforcement learning (DRL) could support agile and adaptive C2 in multi-domain forces. This is particularly relevant for commanders and staff aiming to exploit rapidly evolving situations and fleeting windows of superiority on the battlefield. COA-GPT embodies this concept of agility and adaptability, leveraging LLMs to rapidly develop and modify COAs in response to evolving battlefield conditions. This approach effectively counters the inherent brittleness and extensive training demands typical of DRL methods.
Exploring the synergy between gaming platforms and military training, Goecks et al. (2023) provide insight into how AI algorithms, when combined with gaming and simulation technologies, can be adapted to replicate aspects of military missions [25]. Similarly, COA-GPT can be viewed as part of this trend, utilizing LLMs to enhance the strategic planning phase of military operations and gaming and simulation technologies to evaluate them.
In a similar vein, Waytowich et al. (2022)  [26] demonstrate the application of DRL in commanding multiple heterogeneous actors in a simulated command and control task, modeled on a military scenario within StarCraft II. Their findings indicate that agents trained via an automatically generated curriculum can match or even surpass the performance of human experts and state-of-the-art DRL baselines. This approach to training AI systems in complex scenarios aligns with the objectives of COA-GPT, which seeks to combine human expertise with AI efficiency in generating military COAs, and it is used here as the main baseline comparison for the proposed method, COA-GPT.
In addition to these developments, Schwartz (2020) [27] delves into the application of AI in the Army’s MDMP, specifically in the COA Analysis phase. This research is particularly relevant to COA-GPT, as it demonstrates how AI can assist commanders and their staff in quickly developing and optimizing multiple courses of action in response to the complexities of modern, hyper-contested battlefields. The study also highlights the increasing importance of Multi-Domain Operations (MDO) and the challenges presented by near-peer adversaries equipped with advanced Anti-Access Area Denial (A2AD) capabilities. The COA-GPT aligns with this vision by providing a tool to enhance the MDMP, particularly in the rapid generation and analysis of COAs.
In contrast to these AI-driven approaches, traditional military planning processes, as outlined in U.S. Army’s ATP 5-0.2 [28], involve comprehensive guidelines for staff members in large-scale combat operations. While providing a consolidated source of key planning tools and techniques, these traditional methodologies often lack the agility and adaptability offered by modern AI systems in rapidly changing combat environments. COA-GPT represents a paradigm shift from these traditional methods, bringing the speed and flexibility of LLMs to military planning processes.
IV Methods
@@
COA-GPT is prompted to understand that it serves as a military commander’s assistant to aid C2 personnel in developing COAs. It is aware that its inputs will include mission objectives, terrain information, and details on friendly and threat forces as provided by the commander in text and/or image format. Furthermore, COA-GPT has access to a condensed version of military doctrine, covering forms of maneuver (envelopment, flank attack, frontal attack, infiltration, penetration, and turning movement), offensive tasks (movement to contact, attack, exploitation, and pursuit), and defensive tasks (area defense, mobile defense, and retrograde).
As depicted in Figure 1, the COA-GPT assistant communicates with C2 personnel via natural language. It receives mission-related information such as objectives, terrain details, friendly and threat force description and arrangement, and any planning assumptions the C2 staff might have. For the LLM in the backed, we use OpenAI’s GPT-4-Turbo (named “gpt-4-1106-preview” in their API) for text-only experiments and GPT-4-Vision (named “gpt-4-vision-preview”) for tasks where mission information is given in both text and image format.
@@
V Experiments
V-A Scenario and Experimental Setup
@@
In PySC2, the scenario is realized by mapping StarCraft II units to their military equivalents, adjusting attributes like weapon range, damage, unit speed, and health. For instance, M1A2 Abrams combat armor units are represented by modified Siege Tanks in tank mode, mechanized infantry by Hellions, among others [24]. The Friendly Force consists of 9 Armor, 3 Mechanized infantry, 1 Mortar, 2 Aviation, and 1 Reconnaissance unit. The Threat Force includes 12 Mechanized infantry, 1 Aviation, 2 Artillery, 1 Anti-Armor, and 1 Infantry unit. A specially designed StarCraft II map [24] depicts the TigerClaw area of operations, as shown in Figure 3.
@@ V-B COA Generation
Additionally, for the experiments using LLM with image processing capabilities, COA-GPT takes as input a frame of a Common Operational Picture (COP) that overlays force arrangements in a satellite image of the terrain, as depicted in Figure 4.
@@
Additionally, COA-GPT is programmed with knowledge of specific game engine functions to command each asset, such as:
@@
For integration with the PySC2 game engine, COA-GPT generates COAs in JSON format. These are subsequently translated into function calls understood by the PySC2 engine. A complete COA includes a mission name, strategy description, and specific commands for each asset. Within each simulation rollout, each asset adheres to its assigned command throughout the game, with COA-GPT limited to issuing only one command per asset at the beginning of the simulation.
V-C Human Feedback
@@
To ensure a consistent and fair comparison across all generated COAs, we have standardized the human feedback process. The specific instructions provided in each feedback iteration for the COAs that incorporated human input are as follows:
  • First Iteration: “Make sure both our aviation units directly engages the enemy aviation unit.”
@@
V-D Evaluation Metrics
The evaluation of the generated COAs is based on two key metrics recorded during the COA analysis: total reward and casualty figures. These metrics include:
@@ V-D Evaluation Metrics
V-E Baselines
In this study, we benchmark our proposed method against the two best-performing published approaches for the task: Autocurriculum Reinforcement Learning from a Single Human Demonstration [26] and the Asynchronous Advantage Actor-Critic (A3C) algorithm in Reinforcement Learning [30, 26]. Our comparison includes:
@@ V-E Baselines
  • Reinforcement Learning. As detailed in [26], this baseline employs the A3C algorithm, using either images or vector representations as inputs. Unlike the previous method, it does not incorporate a guiding curriculum for the learning agent. Both input scenarios are evaluated against our method, labeled as “RL-Vec” and “RL-Im”, respectively, in our results.
  • COA-GPT. This serves as an ablation study of our proposed approach, as detailed in Section IV, but without human feedback. Labeled “COA-GPT” (text-only inputs) and “COA-GPT-V” (for experiments with vision models, text and image inputs), this version generates COAs based solely on the initial mission data provided by the C2 personnel, without further human input.
  • COA-GPT with Human Feedback. This is our fully realized method, outlined in Section IV. Beyond the initial mission information from the C2 personnel, this version also incorporates the impact of human feedback on the COA performance. We assess the changes after the first iteration of feedback, denoted as “COA-GPT+H1”, and after the second iteration, labeled “COA-GPT+H2” in the results. Similarly, for experiments using the LLM with vision capabilities the experiments are labelled “COA-GPT-V+H1” and “COA-GPT-V+H2”.
@@
V-F Results
For a comprehensive evaluation of COA-GPT, we generated five COAs for each method variant and conducted ten simulations for each, totaling 50 evaluation rollouts per baseline. All results discussed in this section represent the mean and standard deviation calculated across these 50 rollouts for each evaluation metric. The data for the Reinforcement Learning and Autocurriculum Reinforcement Learning baselines are sourced directly from their respective published work [26].

### 28 Mar 2024 v2

@@ Abstract
The development of Courses of Action (COAs) in military operations is traditionally a time-consuming and intricate process. Addressing this challenge, this study introduces COA-GPT, a novel algorithm employing Large Language Models (LLMs) for rapid and efficient generation of valid COAs. COA-GPT incorporates military doctrine excerpts and domain expertise to LLMs through in-context learning, allowing commanders to input mission information – in both text and image formats – and receive strategically aligned COAs for review and approval. Uniquely, COA-GPT not only accelerates COA development, producing initial COAs within seconds, but also facilitates real-time refinement based on commander feedback. This work evaluates COA-GPT in a military-relevant scenario within a militarized version of the StarCraft II game, comparing its performance against an expert human and state-of-the-art reinforcement learning algorithms. Our results demonstrate COA-GPT’s superiority in generating strategically sound COAs more swiftly, with the added benefits of enhanced adaptability and alignment with commander intentions. COA-GPT’s capability to rapidly adapt and update COAs during missions presents a transformative potential for military planning, particularly in addressing planning discrepancies and capitalizing on emergent windows of opportunity. Performance videos of our method can be seen at https://sites.google.com/view/coa-gpt.
@@ I Introduction
The development of Courses of Action (COAs), a core element of decision-making, highlights these challenges. Traditionally, COA development is meticulous and time-intensive, relying heavily on military expertise. The demands of modern warfare require more efficient COA development and analysis methods.
@@
COA-GPT leverages LLMs to swiftly develop valid COAs, integrating military doctrine excerpts and domain expertise directly into the system’s initial prompts. Commanders can input mission specifics and force descriptions – in both text and image formats - receiving multiple, strategically aligned COAs in a matter of seconds.
@@
  • COA-GPT, a novel framework leveraging LLMs for accelerated development and analysis of COAs, integrating military doctrine excerpts and domain expertise effectively.
@@ II Military Relevance
The current C2 processes in military operations are predominantly linear and involve numerous sequential steps, which can be cumbersome and slow in the fast-paced environment of modern warfare [3, 2]. The advent of AI in military planning, particularly the COA-GPT system, presents an opportunity to radically streamline these processes.
COA-GPT significantly enhances the MDMP by enabling the concurrent development and analysis of COAs. This facilitates quicker decision-making and enhances operation control, leading to the swift generation of actionable intelligence for informed command decisions. Integrating AI tools like COA-GPT in C2 processes can reduce military operations’ physical footprint, lessening the need for extensive personnel and logistics, and aligning with the vision of distributed command structures.
Solutions such as COA-GPT can integrate with war gaming simulators and battlefield sensor data in real-time, enabling quick adaptation to battlefield dynamics. This approach emphasizes early COA analysis integration for swift optimization and comparison [5]. Additionally, by integrating human feedback, COA-GPT allows personnel to adapt AI-generated COAs with their expertise and situational awareness for informed final selections. C2 personnel dynamically interact with COA-GPT to specify objectives, input data, set planning constraints, and adjust proposed COAs, ensuring decisions align with strategic intent and situational requirements.
@@ III-A Planning with Large Language Models
The integration of Large Language Models (LLMs) in the plan of action development has been applied to various sectors, including disaster response operations. Similar to military operations, disaster response demands rapid, informed decision-making under severe time constraints and high-pressure conditions [15, 16, 17]. Traditionally, action plan development in such contexts has been a laborious process, heavily reliant on the experience and expertise of the personnel involved. Given the stakes involved, where delays can result in loss of life, there is a critical need for more efficient and reliable plan development methodologies [18, 19, 15, 20]. In the realm of disaster response, the DisasterResponseGPT [21] algorithm leverages the capabilities of LLMs to quickly generate viable action plans, integrating essential disaster response guidelines within the initial prompt.
The recent studies by Ahn et al. (2022) [12] and Mees et al. (2023) [14] expand the scope of LLM applications to include robotic and visual affordances, illustrating how LLMs can be grounded in real-world contexts. Continuing to show the potential of LLMs for planning, the introduction of Voyager by Wang et al. (2023) [22], an LLM-powered embodied lifelong learning agent in Minecraft, represents an important step in autonomous, continual learning and skill acquisition in complex, ever-changing environments. Similarly, the STEVE-1 [23] model showcases the potential of LLMs in guiding agents to follow complex instructions in virtual settings, leveraging advancements in text-conditioned image generation and decision-making algorithms.
The integration of LLMs into planning has significantly advanced rapid decision-making capabilities in a diverse range of scenarios. COA-GPT builds upon these foundations and demonstrates the versatility of LLMs in translating strategic concepts into actionable plans for military operations.
@@ III-B AI for Military Planning and Operations
The application of AI and LLMs in military planning and operations is a field of growing interest and significant potential. The First-Year Report of ARL Director’s Strategic Initiative, focusing on Artificial Intelligence for Command and Control (C2) of Multi-Domain Operations [24], exemplifies this trend. The report discusses ongoing research into whether AI could support agile and adaptive C2 in multi-domain forces.
Exploring the synergy between gaming platforms and military training, Goecks et al. (2023) provide insight into how AI algorithms, when combined with gaming and simulation technologies, can be adapted to replicate aspects of military missions [25]. In a similar vein, Waytowich et al. (2022)  [26] demonstrate the application of deep reinforcement learning (DRL) in commanding multiple heterogeneous actors in a simulated command and control task, modeled on a military scenario within StarCraft II. Their findings indicate that agents trained via an automatically generated curriculum can match or even surpass the performance of human experts and state-of-the-art DRL baselines.
In addition to these developments, Schwartz (2020) [27] delves into the application of AI in the Army’s MDMP, specifically in the COA Analysis phase. They demonstrate how AI can assist commanders and their staff in quickly developing and optimizing multiple courses of action in response to the complexities of modern, hyper-contested battlefields. The study also highlights the increasing importance of Multi-Domain Operations (MDO) and the challenges presented by near-peer adversaries equipped with advanced Anti-Access Area Denial (A2AD) capabilities.
In contrast to these AI-driven approaches, traditional military planning processes, as outlined in U.S. Army’s ATP 5-0.2 [28], involve comprehensive guidelines for staff members in large-scale combat operations. While providing a consolidated source of key planning tools and techniques, these traditional methodologies often lack the agility and adaptability offered by modern AI systems in rapidly changing combat environments.
In the evolving landscape of military planning and operations, the integration of AI and LLMs signifies a shift towards more agile and adaptive command and control strategies. COA-GPT exemplifies this shift, leveraging LLMs for accelerated COA development, thus addressing the limitations of manual planning processes and previous automated approaches.
IV Methods and Experiments
@@
COA-GPT is prompted to understand that it serves as a military commander’s assistant to aid C2 personnel in developing COAs. It is aware that its inputs will include mission objectives, terrain information, and details on friendly and threat forces as provided by the commander in text and/or image format. It is also instructed to use specific commands to assign tasks for each asset in the friendly forces. Furthermore, COA-GPT has access to military doctrine excerpts, covering forms of maneuver (envelopment, flank attack, frontal attack, infiltration, penetration, and turning movement), offensive tasks (movement to contact, attack, exploitation, and pursuit), and defensive tasks (area defense, mobile defense, and retrograde). The complete system prompt and doctrinal excerpts given to COA-GPT are reproduced in Appendix -A.
As depicted in Figure 1, the COA-GPT assistant communicates with C2 personnel via natural language. It receives mission-related information such as objectives, terrain details, friendly and threat force description and arrangement, and any planning assumptions the C2 staff might have. For the LLM in the back end, we use OpenAI’s GPT-4-Turbo (named “gpt-4-1106-preview” in their API) for text-only experiments and GPT-4-Vision (named “gpt-4-vision-preview”) for tasks where mission information is given in both text and image format.
@@
IV-A Scenario and Experimental Setup
@@
In PySC2, the scenario is realized by mapping StarCraft II units to their military equivalents, adjusting attributes like weapon range, damage, unit speed, and health. For instance, M1A2 Abrams combat armor units are represented by modified Siege Tanks in tank mode, mechanized infantry by Hellions, among others [24] (see Appendix -C for more details). The Friendly Force consists of 9 Armor, 3 Mechanized infantry, 1 Mortar, 2 Aviation, and 1 Reconnaissance unit. The Threat Force includes 12 Mechanized infantry, 1 Aviation, 2 Artillery, 1 Anti-Armor, and 1 Infantry unit.
In terms of terrain, a custom StarCraft II map [24] was made to depict the TigerClaw area of operations, as shown in Figure 3. In this scenario, the Threat Force is controlled by the default in-game StarCraft II AI. They are programmed to automatically move to fixed defensive positions to defend the three main Thar Thar Wadi crossings. Upon detecting the Friendly Force within range, they open fire. Additionally, the Threat Aviation unit performs a fixed patrol route covering the three crossing points, while also engaging any Friendly Force assets that come into range.
@@ IV-B COA Generation
For the experiments using LLM with image processing capabilities, COA-GPT takes as input a frame of a Common Operational Picture (COP) that overlays force arrangements in a satellite image of the terrain, as depicted in Figure 4.
@@
Mission objective, terrain, and force information are integrated into a single COA generation prompt, as shown in Appendix -B, that is then used to query the LLM in the back end.
For integration with the PySC2 game engine, COA-GPT outputs COAs in JSON format, as detailed in the example COA statement found in Appendix -A. These task allocations from COA-GPT are translated into specific function calls compatible with the PySC2 engine [^1] PySC2 Action Definition: https://github.com/google-deepmind/pysc2/blob/master/pysc2/lib/actions.py., facilitating in-game actions based on the generated COAs. COA-GPT is programmed with knowledge of the specific game engine functions to command each asset:
@@
A comprehensive COA encompasses a mission name, a brief strategic description, and designated commands for each asset involved, as defined by the presented game engine functions. During a simulation rollout, each asset strictly follows its initial command issued by COA-GPT at the start of the simulation. For instance, a command to $engage\_target\_unit(X,Y,coord_{x},coord_{y})$ would specify unit $X$ to engage target unit $Y$, maintaining this engagement until the battle concludes or a new COA is issued. Similarly, an $attack\_move\_unit(Z,coord_{x},coord_{y})$ command directs unit $Z$ to move to a specified map coordinate, engaging any hostile units encountered en route. The PySC2 engine processes these commands at every simulation time step to ensure each asset is correctly controlled.
IV-C Human Feedback
@@
To ensure a consistent and fair comparison and evaluation process across all generated COAs, we have standardized the human feedback process specifically for this work. The specific instructions provided in each feedback iteration for the COAs that incorporated human input are as follows:
  • First Iteration: “Make sure both our aviation units directly engage the enemy aviation unit.”
@@
IV-D Evaluation Metrics
The evaluation of the generated COAs is based on three key metrics recorded during the COA analysis: total reward, friendly force casualties, and threat force casualties:
@@ IV-D Evaluation Metrics
IV-E Baselines
In this study, we benchmark our proposed method against an expert human and the two best-performing published approaches for this scenario: Autocurriculum Reinforcement Learning from a Single Human Demonstration [26] and the Asynchronous Advantage Actor-Critic (A3C) algorithm in Reinforcement Learning [30, 26]:
@@ IV-E Baselines
  • Reinforcement Learning. As detailed in [26], this baseline employs the A3C algorithm, using either images or vector representations as inputs. Unlike the previous method, it does not incorporate a guiding curriculum for the learning agent. Both input modes are evaluated against our method, labeled as “RL-Vec” and “RL-Im”, respectively, in our results.
  • COA-GPT. This serves as an ablation study of our proposed approach but without human feedback. Labeled “COA-GPT” (text-only inputs) and “COA-GPT-V” (for experiments with vision models using both text and image inputs), this version generates COAs based solely on the initial mission data provided by the C2 personnel, without further human input.
  • COA-GPT with Human Feedback. This is our fully realized method. Beyond the initial mission information from the C2 personnel, this version also incorporates the impact of human feedback on the COA performance. We assess the changes after the first iteration of feedback, denoted as “COA-GPT+H1”, and after the second iteration, labeled “COA-GPT+H2” in the results. Similarly, for experiments using the LLM with vision capabilities the experiments are labeled “COA-GPT-V+H1” and “COA-GPT-V+H2”.
  • Expert Human. A human expert was tasked to complete the same scenario over 15 trials. Differently from COA-GPT, the human controlled the assets using a mouse a keyboard in real-time throughout the whole game. Similarly, the human user was only able to command attack-move or direct engage commands.
@@
V Experimental Results
V-A Qualitative Results
To illustrate the full realization of the proposed methodology, Figure 6 depicts the evolution of a generated COA in response to human feedback. Initially, the COA generated without human input (Figure 6a) displays a certain configuration for the planned movement of friendly forces across bridges and their engagement with hostile units. After the first round of human feedback (“Make sure both our aviation units directly engage the enemy aviation unit.”), we see a strategic pivot (Figure 6b); the friendly aviation units are now tasked to engage the enemy’s aviation assets directly. This adjustment reflects the human commander’s intent to prioritize air superiority. The second iteration of feedback — “Make sure only our Scout unit is commanded to control Bridge Bobcat (x: 75 y: 26) and our other assets (not the aviation) are split into two groups and commanded to move towards both enemy artillery using the attack_move command.” — as seen in Figure 6c), results in a more nuanced approach: the friendly forces are divided, with specific units tasked to target enemy artillery positions. Additionally, the reconnaissance unit is ordered to secure the northern bridge and the friendly aviation is still tasked to engage the threat aviation, demonstrating that COA-GPT successfully followed the commander’s intent conveyed via textual feedback.
V-B Quantitative Results
For a comprehensive evaluation of COA-GPT, we generated five COAs for each method variant and conducted ten simulations for each, totaling 50 evaluation rollouts per baseline. All results discussed in this section represent the mean and standard deviation calculated across these 50 rollouts for each evaluation metric. It is important to note that the game of StarCraft II is mostly deterministic [29] - with stochasticity mostly being injected via weapon speed in the case of our TigerClaw scenario - which implies that the variance present in the results originates from the diversity of the COAs generated by COA-GPT. The data for the Reinforcement Learning and Autocurriculum Reinforcement Learning baselines are sourced directly from their respective published work [26]. Performance videos of COA-GPT can be seen at the project webpage  [^2] COA-GPT Project Webpage: https://sites.google.com/view/coa-gpt..